### v1

@@ 2.4 Inverse design and parametric learning
Oftentimes, in science and engineering, researchers are not only interested in the forward solution of a particular problem, but are keen to achieve the solutions of suitably defined inverse problems. That is, they are interested in the geometric properties or in the determination of fundamental parameters of the model that are consistent with a desired performance. This class of problems, known as inverse problems, are often ill-posed (e.g., they do not have a unique solution) and are difficult and time-consuming to solve using traditional methods [33]. In particular, relevant inverse problems feature non-convex objective functions leading to a very complex solution landscape with multiple local minima, making it impossible to find global optimal solutions. For these reasons, they have traditionally been approached by constrained optimization methods that, through numerous iterations of a forward model, reconstruct simultaneously both the target function and the relevant physical properties of the system [22, 34]. Therefore, inverse problems are computationally more expensive than the corresponding forward problems due to the high dimensional nature of the parameter space of solutions that the solver needs to navigate in order to approach an optimal one. Despite their computational challenges, inverse problems have ubiquitous applications in science and engineering, ranging from MRI technologies to geophysics and remote sensing. However, in the context of semiconductor devices, a systematic approach to tackle relevant inverse problems is still missing, despite some recent attempts at extracting functional parameters using ML-based methods by Ling-Feng Mao et al. and Djordjevic et al. [35, 36]. Developing a systematic approach to solve inverse problems for semiconductor devices could enable a faster exploration of optimal device designs and properties. In addition, a realistic inverse design framework should be flexible enough to incorporate fabrication and material constraints effortlessly, thus enabling inverse design capabilities.
The 1D and 2D problems described thus far showed that DDNet can solve traditional forward semiconductor problems with precision comparable to state-of-the-art numerical solvers. However, DDNet is not intended to replace traditional, well-established forward numerical solvers, but rather to complement them by providing support for applications where they encounter difficulties. For this reason, we show in Fig. 4 the range of inverse solutions that have been implemented and validated using DDNet so far. In Fig. 4a, we display the modified DDNet architecture developed to solve inverse problems. First, by directly encoding relevant device parameters within the input training vector of DDNet, here schematically represented by $P_{1},\dots,P_{n}$ in panel a, we achieve parametric learning and robust extrapolation of device properties. For example, in the case of a $p$-$i$-$n$ diode geometry, used extensively as high voltage rectifiers or avalanche photodetectors, we set up a training simulation where the width of the intrinsic region was incorporated as an additional learning parameter to DDNet, along with its spatial dimension. By simultaneously solving the drift-diffusion equations for four different doping conditions, we could then extrapolate the functional response of the electric field within the device region, which has been plotted against the width of the intrinsic region in panel b.
@@
In addition, we show in panels c and d examples of performance-based inverse design discovery. In the first example, we aim at finding the slope of the doping profile $C(x)$ for a $p$-$n$ junction that features a desired electric field distribution specified at several user-defined points inside the device domain. This setup corresponds to a genuine inverse design problem, as opposed to simpler parameter retrieval problems [17]. To solve inverse problems, we modify the $n$-Net and $p$-Net subnetworks to include the additional outputs $\hat{N}_{d}^{+}$ and $\hat{N}_{a}^{-}$, respectively, which specify the doping concentration of the inversely retrieved device structure $\hat{C}(\boldsymbol{x})$. Furthermore, we replace Dirichlet conditions on $\phi$ with a small set of target electric field data points, which are shown with red circles in panel a. An additional inverse loss term is added, allowing DDNet to learn both the doping profile and the forward solution simultaneously. Results for this first example are shown in panel c: the recovered doping profile has the appropriate slope to reproduce the desired target performance, while the predicted electric potential and carrier densities remain accurate. This example shows that, despite the ill-posed nature of this inverse problem, DDNet recovers a realistic doping profile from the family of possible solutions. As a second example, we attempt to inversely design an unknown device configuration (i.e., doping profile) following the same principle, as shown in panel d. This time we select $N=5$ points that we would like to match to a different electric field distribution on the device domain. However, we might be unaware of the existence of a device with such features. Using the same DDNet architecture, we recover a doping profile that exhibits the desired electric field distribution, as shown by the results in panel d. A posteriori, we would learn that this structure represents a BJT-like (or $p$-$n$-$p$ junction) device with symmetric doping. However, the solution of the inverse design problem was guided solely by the target characteristics, not the material geometry, under the constraint of the total length. Rather, the material properties, i.e., the doping profile, were derived during the training of DDNet to satisfy the drift-diffusion equations. For both examples c and d, we have overlaid the position of the target data points (red circles) on the reconstructed forward simulation of the electric potential.
Furthermore, we demonstrate the capabilities of DDNet in relation to two analytically intractable problems of inverse semiconductor device design. First, we attempt the retrieval of the doping profile for a $1\mu\text{m}$ avalanche photodiode (APD), characterized by a high electric field response at equilibrium, which under reverse bias enables the amplification of small photogenerated signals. The set-up of this inverse design problem is presented in Fig. 5a, where the target electric field response is plotted over the device domain. After 3 minutes of training time, DDNet performs a full inversion of the doping profile of an APD, as shown in panel b. Interestingly, DDNet retrieves the $25$ nm thick $n^{+}$ doped region needed to enhance the electric field response after the avalanche region between the intrinsic ($i$) and $p^{+}$ doped regions, as shown schematically in b. This doping inversion was performed only with prior knowledge of the desired electric response of the device subject to the drift-diffusion equations, and DDNet’s training required no additional time or computational costs compared to the corresponding 1D forward simulations.
@@
Figure 5: RF sensing and fabrication-tailored inverse design applications of DDNet. a, Multiscale electric field target performance dataset for a prototypical RF sensing device. b, DDNet’s inverted doping profile with a comparison to a true APD doping and corresponding APD device schematic. c, Target electric potential inversion dataset. d, Binarization retrieval process during DDNet’s inverse design, where an impractical grated doping profile (left) is slowly binarized during the process until a binarized one is obtained (right). e, Forward simulation with the binary doping profile shown in d using TCAD Sentaurus, in close agreement with the target.
@@
As our last example, we demonstrate how to include fabrication constraints during the inverse design process. In Fig. 5c we show the two-dimensional plot of the electric potential from the junction discussed in 2.3. This electric potential profile was obtained using the continuous doping profile shown in Fig. 3b. However, we now want to inversely obtain a binarized doping profile that produces a device performance as close as possible to the original (non-binarized) situation. To achieve this, we set up our full inverse design of the doping profile by providing DDNet the 2D data of the target electric potential, and only specifying the values of $\hat{p}$ and $\hat{n}$ at the source and drain contacts. A first solution, which is obtained with no additional constraints imposed on $\hat{C}$, is shown in the first figure of panel d. This figure shows a graded doping profile that, despite satisfying the Poisson equation, has limited applicability, since it cannot be practically implemented. To avoid retrieving impractical solutions, we borrow an inverse design idea from topological optimization [37], which has recently been employed in the inverse design of photonic structures [23]. Namely, we slowly binarize the DDNet output $\hat{C}$ during the training process with a custom thresholding function $\sigma_{cst}(\boldsymbol{y},t)$ to ensure that the output does not deviate from our fabrication requirements. The result is shown on the right of panel d, where a fully binary doping profile has been obtained. To validate the accuracy of the newly retrieved structure, we perform a forward simulation in TCAD Sentaurus and plot $\phi$ in Fig. 5e. The relative $L_{1}$ error between the target and TCAD $\phi$ was lower than 1%. Similar to the 1D case, the 2D inversion problem added no computational overhead compared to the forward problem. These inverse design examples highlight one of the key strengths of DDNet: the ability to solve inverse PDE problems without incurring additional computational cost or requiring large datasets.
@@ 3 Discussion
We introduced DDNet, a unified physics-informed deep learning framework to solve forward, parametric, and inverse design problems in semiconductor engineering. DDNet trains on synthetic data, but has the infrastructure to augment its predictive power using experimental or target performance datasets for inverse design applications with minimal computational overhead. Furthermore, the architecture shown here can be fine-tuned to solve semiconductor problems for different materials and geometries, including problems in three spatial dimensions and heterostructures. The novel framework presented complements and expands the capabilities of traditional forward solvers and data-driven machine learning algorithms, demonstrating inverse design capabilities and functionalities not explored before. Future research will focus on scaling and generalizing DDNet to more complex device geometries, multi-material interfaces, and coupled multi-physics applications, such as electrical and thermal transport.
@@ 4 Methods
4.1 Boundary conditions of the drift-diffusion equations
For completeness, we rewrite the steady-state drift-diffusion equations:
$\epsilon\Delta\phi$ $=q(n-p-C),$
$\nabla\cdot\boldsymbol{J}_{n}$ $=qR,$
$\nabla\cdot\boldsymbol{J}_{p}$ $=-qR,$
$\boldsymbol{J}_{n}$ $=q\mu_{n}(U_{t}\nabla n-n\nabla\phi),$
$\boldsymbol{J}_{p}$ $=-q\mu_{p}(U_{t}\nabla p+p\nabla\phi)$ (2)
@@ 4.1 Boundary conditions of the drift-diffusion equations
$\phi(x_{c})$ $=U_{t}\ln\left(\frac{n}{n_{i}}\right)-V_{a},\quad\text{for }n\text{-type contact},$ (3)
$\phi(x_{c})$ $=-U_{t}\ln\left(\frac{p}{n_{i}}\right),\quad\text{for }p\text{-type contact},$ (4)
@@
where $V_{\text{a}}$ is the externally applied potential at the contact, $n_{i}$ is the intrinsic carrier concentration, and $U_{t}=kT/q$ is the thermal voltage. These expressions follow from the condition of charge neutrality np=Cn-p=C, and the Boltzmann approximation $n=n_{i}\exp(\phi/U_{t})$, $p=n_{i}\exp(-\phi/U_{t})$.
@@
4.2 Logarithmic DDNet formulation
The modified drift-diffusion equations used to train DDNet are obtained by rescaling Eq. 1 according to the scaling factors in [6, Table 2.4.1]:
@@ 4.2 Logarithmic DDNet formulation
In Eq. 10, we have introduced the scaling constant $\lambda=\lambda_{D}/\ell$, where $\lambda_{D}=\sqrt{\frac{\epsilon_{s}U_{t}}{q\tilde{C}(x)}}$ is the Debye length of the device, and $\ell$ is a characteristic device length chosen such that it is of the same order of magnitude as the diameter of the semiconductor simulation domain $\Omega$, $\ell=O(\text{diameter}(\Omega))$. A complete description of the scalings and variables can be found in the Supplementary Information. The equations in 10 are obtained to aid with the numerical treatment of the functions $\hat{\phi}$, $\hat{n}$, and $\hat{p}$ [6]. In particular, this formulation helps dealing with the large dynamic ranges of $\hat{n}$ and $\hat{p}$, which typically span between twelve and sixteen orders of magnitude for silicon, but they can reach up to forty orders of magnitude for materials like GaN. To do so, $\hat{n}$ and $\hat{p}$ are rescaled by $\tilde{C}$, defined as $\tilde{C}=O(C_{\text{max}})$, which ensures that the maximum value of $\hat{n}$ and $\hat{p}$ is unity, while $\hat{\phi}$ is rescaled by $U_{t}$.
@@
To capture the full dynamic ranges of $\hat{n}$ and $\hat{p}$ we rescale the outputs according to $\hat{p}/\tilde{C}$ and $\hat{n}/\tilde{C}$ and train according to the newly defined variables -log($\hat{p}/\tilde{C}$) or -log($\hat{n}/\tilde{C}$), which compress their dynamic range. Finally, the outputs are exponentially transformed through a hard constraint [31] as $e^{-\text{log}(\hat{n}/\tilde{C})}$ and $e^{-\text{log}(\hat{p})}$, thereby recovering the full dynamic range of the solutions. For a more careful discussion of this logarithmic formulation and a comparison with a traditional architecture, we refer the reader to the Supplementary Information.
4.3 DDNet training
[…]
Each loss term $\mathcal{L}_{t}$ is computed using a mean squared error formula to minimize the residual error on $\mathcal{N}_{i}$ collocation or boundary points using the ADAM optimizer [38]:
[…]
@@ 4.3 DDNet training
After the optimization step is completed, the network parameters $\tilde{\theta}$ are updated using backpropagation, and the process is repeated until a threshold error is achieved, or the training routine has reached a set number of epochs. To train DDNet, we typically use $2^{12}=4096$ uniform random collocation points on the PDE domain and $2^{8}=256$ uniform random collocation points on the boundary for a total of $N=40,000$ epochs. Additionally, we introduce a resampling routine during training that re-generates the collocation points every $N/f$ epochs, where $f$ is an adjustable frequency set to 100 by default. We have found this technique to improve with convergence and precision without introducing a significant additional overhead. All the DDNet codes and training routines have been developed in-house using the TensorFlow [39] machine learning package, and the scripts ran on a single L40S GPU of the Boston University Shared Supercomputing Cluster (SCC). We have included a comparison of runtime on different GPUs in the Supplementary Information.
[…]
@@
Supplementary information We have included additional data on semiconductors theory background, DDNet training and network architecture comparisons, and convergence times and loss functions in the Supplementary Information.
[…]
@@
Acknowledgements This work was supported by the U.S. Army Research Office, RF-Center managed by Dr. T. Oder (Grant #W911NF-22-2-0158).

### v2

@@ 2.4.1 DDNet for inverse problems and parametric learning
Oftentimes, in science and engineering, researchers are not only interested in the forward solution of a particular problem, but are keen to achieve the solutions of suitably defined inverse problems. That is, they are interested in the geometric properties or in the determination of fundamental parameters of the model that are consistent with a desired performance. This class of problems, known as inverse problems, are often ill-posed (e.g., they do not have a unique solution) and are difficult and time-consuming to solve using traditional methods [burger_inverse_2004]. In particular, relevant inverse problems feature non-convex objective functions leading to a very complex solution landscape with multiple local minima, making it impossible to find global optimal solutions. For these reasons, they have traditionally been approached by constrained optimization methods that, through numerous iterations of a forward model, reconstruct simultaneously both the target function and the relevant physical properties of the system [chen_computational_2018, colton_inverse_2019]. Therefore, inverse problems are computationally more expensive than the corresponding forward problems due to the high dimensional nature of the parameter space of solutions that the solver needs to navigate in order to approach an optimal one. Despite their computational challenges, inverse problems have ubiquitous applications in science and engineering, ranging from MRI technologies to geophysics and remote sensing. However, in the context of semiconductor devices, a systematic approach to tackle relevant inverse problems is still missing, despite some recent attempts at extracting functional parameters using ML-based methods by Ling-Feng Mao et al. and Djordjevic et al. [djordjevic_inverse_2021, mao_addressing_2024]. Developing a systematic approach to solve inverse problems for semiconductor devices could enable a faster exploration of optimal device designs and properties. In addition, a realistic inverse design framework should be flexible enough to incorporate fabrication and material constraints effortlessly, thus enabling inverse design capabilities. The 1D and 2D problems described thus far showed that DDNet can solve traditional forward semiconductor problems with precision comparable to state-of-the-art numerical solvers. However, DDNet is not intended to replace traditional, well-established forward numerical solvers, but rather to complement them by providing support for applications where they encounter difficulties. For this reason, we show in Fig. 4 the range of inverse solutions that have been implemented and validated using DDNet so far. In Fig. 4a, we display the modified DDNet architecture developed to solve inverse problems. First, by directly encoding relevant device parameters within the input training vector of DDNet, here schematically represented by $P_{1},\dots,P_{n}$ in panel a, we achieve parametric learning and robust extrapolation of device properties. For example, in the case of a $p$-$i$-$n$ diode geometry, used extensively as photodetectors, we set up a training simulation where the width of the intrinsic region was incorporated as an additional learning parameter to DDNet, along with its spatial dimension. By simultaneously solving the drift-diffusion equations for four different doping conditions, we could then extrapolate the functional response of the electric field within the device region, which has been plotted against the width of the intrinsic region in panel b. In addition, we show in panels c and d examples of performance-based inverse design discovery. In the first example, we aim at finding the slope of the doping profile $C(x)$ for a $p$-$n$ junction that features a desired electric field distribution specified at several user-defined points inside the device domain. This setup corresponds to a genuine inverse design problem, as opposed to simpler parameter retrieval problems [riganti_auxiliary_2023]. To solve inverse problems, we modify the $n$-Net and $p$-Net subnetworks to include the additional outputs $\hat{N}_{d}^{+}$ and $\hat{N}_{a}^{-}$, respectively, which specify the doping concentration of the inversely retrieved device structure $\hat{C}(\boldsymbol{x})$.
@@
Furthermore, we replace Dirichlet conditions on $\phi$ with a small set of target electric field data points, which are shown with red circles in panel a. An additional inverse loss term is added, allowing DDNet to learn both the doping profile and the forward solution simultaneously. Results for this first example are shown in panel c: the recovered doping profile has the appropriate slope to reproduce the desired target performance, while the predicted electric potential and carrier densities remain accurate. This example shows that, despite the ill-posed nature of this inverse problem, DDNet recovers a realistic doping profile from the family of possible solutions.
As a second example, we attempt to inversely design an unknown device configuration (i.e., doping profile) following the same principle, as shown in panel d. This time we select $N=5$ points that we would like to match to a different electric field distribution on the device domain. In addition, we select 3 more points that we want to match with a specified carrier concentration of $n$ and $p$. However, we might be unaware of the existence of a device with such features. Using the same DDNet architecture, we recover a doping profile that exhibits the desired electric field distribution, as shown by the results in panel d. A posteriori, we would learn that this structure represents a BJT-like (or $p$-$n$-$p$ junction) device with symmetric doping. However, the solution of the inverse design problem was guided solely by the target characteristics, not the material geometry, under the constraint of the total length. Rather, the material properties, i.e., the doping profile, were derived during the training of DDNet to satisfy the drift-diffusion equations. For both examples c and d, we have overlaid the position of the target data points (red circles) on the reconstructed forward simulation of the electric potential. Furthermore, we demonstrate the capabilities of DDNet in relation to two analytically intractable problems of inverse semiconductor device design.
@@
Figure 5: Inverse design of APDs and binarized structures using DDNet. a, Multiscale electric field target performance dataset for a prototypical RF sensing device. b, DDNet’s inverted doping profile with a comparison to a true APD doping and corresponding APD device schematic.
2.4.2 Inverse design of an avalanche photodiode (APD)
We first attempt the retrieval of the doping profile for a $1\mu\text{m}$ avalanche photodiode (APD), characterized by a high electric field response at equilibrium, which under reverse bias enables the amplification of small photogenerated signals. The set-up of this inverse design problem is presented in Fig. 5a, where the target electric field response is plotted over the device domain. After 3 minutes of training time, DDNet performs a full inversion of the doping profile of an APD, as shown in panel b. Interestingly, DDNet retrieves the $25$ nm thick $n^{+}$ doped region needed to enhance the electric field response after the avalanche region between the intrinsic ($i$) and $p^{+}$ doped regions, as shown schematically in b. This doping inversion was performed only with prior knowledge of the desired electric response of the device subject to the drift-diffusion equations, and DDNet’s training required no additional time or computational costs compared to the corresponding 1D forward simulations.
Figure 6: Inverse design of a binarized structure using DDNet. a, Target electric potential inversion dataset. b, Binarization retrieval process during DDNet’s inverse design, where an impractical grated doping profile (left) is slowly binarized during the process until a binarized one is obtained (right). c, Forward simulation with the binary doping profile shown in d using TCAD Sentaurus, in close agreement with the target.
2.4.3 Inverse design of binary doping with fabrication constraints
@@
As our last example, we demonstrate how to include fabrication constraints during the inverse design process. In Fig. 6a we show the two-dimensional plot of the electric potential from the junction discussed in 2.3. This electric potential profile was obtained using the continuous doping profile shown in Fig. 3b. However, we now want to inversely obtain a binarized doping profile that produces a device performance as close as possible to the original (non-binarized) situation. To achieve this, we set up our full inverse design of the doping profile by providing DDNet the 2D data of the target electric potential, and only specifying the values of $\hat{p}$ and $\hat{n}$ at the source and drain contacts. A first solution, which is obtained with no additional constraints imposed on $\hat{C}$, is shown in the first figure of panel d. This figure shows a graded doping profile that, despite satisfying the Poisson equation, has limited applicability, since it cannot be practically implemented. To avoid retrieving impractical solutions, we borrow an inverse design idea from topological optimization [christiansen_inverse_2021], which has recently been employed in the inverse design of photonic structures [riganti_multiscale_2025]. Namely, we slowly binarize the DDNet output $\hat{C}$ during the training process with a custom thresholding function $\sigma_{cst}(x,t)=\frac{\gamma}{1+e^{-\xi(t)\cdot x}}+\eta$ to ensure that the output does not deviate from our fabrication requirements. Here, $\sigma_{cst}(x,t)$ is explicitly time-dependent since its slope $\xi(t)$ slowly increases during the training process, allowing for a smooth solution in the early epochs. As the training proceeds, the slope becomes sharper until the doping converges to a perfectly binary solution. The result is shown on the right of panel d, where a fully binary doping profile has been obtained. To validate the accuracy of the newly retrieved structure, we perform a forward simulation in TCAD Sentaurus and plot $\phi$ in Fig. 6c. The relative $L_{1}$ error between the target and TCAD $\phi$ was lower than 1%. Similar to the 1D case, the 2D inversion problem added no computational overhead compared to the forward problem. These inverse design examples highlight one of the key strengths of DDNet: the ability to solve inverse PDE problems without incurring additional computational cost or requiring large datasets.
@@ 3 Discussion
We introduced DDNet, a unified physics-informed deep learning framework to solve forward, parametric, and inverse design problems in semiconductor engineering. As an inverse design tool, DDNet usually trains on synthetic data, but has the infrastructure to augment its predictive power by incorporating experimental or target performance datasets at minimal computational overhead. Furthermore, the architecture shown here can be fine-tuned to solve semiconductor problems for different materials and geometries, including heterostructures. The novel framework presented complements and expands the capabilities of traditional forward solvers and data-driven machine learning algorithms, demonstrating inverse design capabilities and functionalities not explored before. Future research will focus on scaling and generalizing DDNet to more complex device geometries, multiple materials, and coupled multi-physics applications, such as optical and thermal transport.
@@ 4 Methods
4.1 Boundary Conditions
4.1.1 Ohmic Contacts
@@
$\phi(x_{c})$ $=\phi_{0}-V_{a},$ (5)
@@
where $\phi_{0}=U_{t}\ln\left(\frac{n}{n_{i}}\right)$ for n-type contact or $\phi_{0}=-U_{t}\ln\left(\frac{p}{n_{i}}\right)$ for p-type contact. $V_{\text{a}}$ is the externally applied potential at the contact, $n_{i}$ is the intrinsic carrier concentration, and $U_{t}=kT/q$ is the thermal voltage. These expressions follow from the condition of charge neutrality np=Cn-p=C, and the Boltzmann approximation $n=n_{i}\exp(\phi/U_{t})$, $p=n_{i}\exp(-\phi/U_{t})$.
4.1.2 Insulating Boundaries
@@
4.2 Equations rescaling and logarithmic treatment
The modified drift-diffusion equations used to train DDNet are obtained by rescaling Eq. 1 according to the scaling factors in [1985Markowich, Table 2.4.1]:
@@ 4.2 Equations rescaling and logarithmic treatment
In Eq. 11, we have introduced the scaling constant $\lambda=\lambda_{D}/\ell$, where $\lambda_{D}=\sqrt{\frac{\epsilon_{s}U_{t}}{q\tilde{C}(x)}}$ is the Debye length of the device, and $\ell$ is a characteristic device length chosen such that it is of the same order of magnitude as the diameter of the semiconductor simulation domain $\Omega$, $\ell=O(\text{diameter}(\Omega))$. A complete description of the scalings and variables can be found in Section 1 and Table S1 of the Supplementary Information. The equations in 11 are obtained to aid with the numerical treatment of the functions $\hat{\phi}$, $\hat{n}$, and $\hat{p}$ [1985Markowich]. In particular, this formulation helps deal with the large dynamic ranges of $\hat{n}$ and $\hat{p}$, which typically span between twelve and sixteen orders of magnitude for silicon, but they can reach up to forty orders of magnitude for materials like GaN. To do so, $\hat{n}$ and $\hat{p}$ are rescaled by $\tilde{C}$, defined as $\tilde{C}=O(C_{\text{max}})$, which ensures that the maximum value of $\hat{n}$ and $\hat{p}$ is unity, while $\hat{\phi}$ is rescaled by $U_{t}$.
@@
To capture the full dynamic ranges of $\hat{n}$ and $\hat{p}$ we rescale the outputs according to $\hat{p}/\tilde{C}$ and $\hat{n}/\tilde{C}$ and train according to the newly defined variables -log($\hat{p}/\tilde{C}$) or -log($\hat{n}/\tilde{C}$), which compress their dynamic range. Finally, the outputs are exponentially transformed through a hard constraint [lu_physics-informed_2021] as $e^{-\text{log}(\hat{n}/\tilde{C})}$ and $e^{-\text{log}(\hat{p}/\tilde{C})}$, thereby recovering the full dynamic range of the solutions. We have compared the training using the logarithmic scaling with a traditional PINN and included the analysis in Section 3 of the Supplementary Information. We show that the traditional PINN framework achieves precision comparable to DDNet for the electrostatic potential, as shown in Figure S1. However, Figure S2 shows that the traditional PINN fails to capture the full dynamic range for $n$ and $p$.
4.3 Training details
@@
To train DDNet, we typically use $2^{12}=4096$ uniform random collocation points on the PDE domain and $2^{8}=256$ uniform random collocation points on the boundary for a total of $N=40,000$ epochs. Additionally, we introduce a resampling routine during training that re-generates the collocation points every $N/f$ epochs, where $f$ is an adjustable frequency set to 100 by default. We have found this technique to improve with convergence and precision without introducing a significant additional overhead. All the DDNet codes and training routines have been developed in-house using the TensorFlow [tensorflow2015-whitepaper] machine learning package, and the scripts ran on a single L40S GPU of the Boston University Shared Computing Cluster (SCC).
Supplementary information
@@
We have included additional data on DDNet training and examples, network architecture comparisons, convergence times, and loss function convergence in the Supplementary Information.
Funding
@@
This work was supported by the U.S. Army Research Office, RF-Center managed by Dr. T. Oder (Grant #W911NF-22-2-0158).
Author Contributions
E.B. and L.D.N. obtained the grant and initiated this project. L.D.N. conceived the idea and oversaw the technical research. R.R. designed the model, trained the model, and obtained the results. R.R. and L.D.N. wrote the main manuscript with valuable inputs from all the authors. E. B. and M.G.C.A. helped develop relevant application examples. All authors reviewed the manuscript.
Competing Interests
The authors declare no competing interests.